04:- A sand sample is tested in a direct shear device. The vertical normal stress on the sample is 300 kN/m². The horizontal subject / soil mechanics
i want the solution as soon as possible
Answer #1

Solution Given :- verticle normal stress - 300 kN/m² Horizontal shear stress = 210 kn/m2 In direct shear test, As we know tha6 principle stress at failure, a at failure 45t $ 2 Where ø ir angle of internal frictim 62 a = 45°+ 34,992 2 to, 62.4969 Assubstitute the value of eqr ③ in ear ② (6,+)+(-_^)*(-0,5725) 300 300 = (8+62) - 0.5735 X 256.338 6+ 3 = 894.0 equ 6 As we kno(6) liven, principle stress difference = 130 KN 1m2 6- 63 = 130 kN/m2 eq As we know that from ean o 6 = 3.6889 2 6 substitut